# Sublinear Regret for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

### v2

@@ A.2 Mean Increment
$h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ in the updating rule (27). First, note that $\int_{0}^{\cdot}Z_{1,n}^{(2,j)}(t)\mathrm{d}W_{n}^{(j)}(t)$ is a martingale by virtue of Lemma 1 and (33). Taking integration and expectation in (32), we get
$\mathbb{E}[Z_{1,n}(s)]$ $=-\int_{0}^{s}k_{1}(t;\boldsymbol{\theta}_{n})(B+(\sum_{j=1}^{m}C_{j}D_{j})+(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\phi_{1,n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t,$ (34)
@@
$h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ $=-(B+(\sum_{j=1}^{m}C_{j}D_{j})+(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\phi_{1,n})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t$ (35)
@@
$l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})=(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t.$ (36)
Next we study $h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$. It follows from (28) that $a(\phi_{1})$ is a quadratic function of $\phi_{1}$ and
@@
Hence, by (30), we have
$\mathbb{E}[x_{n}(t)^{2}]$ $=(\sum_{j=1}^{m}D_{j}^{\top}\phi_{2,n}D_{j})\int_{0}^{t}e^{a(\phi_{1,n})(t-s)}\mathrm{d}s+x_{0}^{2}e^{a(\phi_{1,n})t}$ (37)
@@
$l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ $=(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t$ (38)
@@
$|h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})|$ $\leq|\phi_{1,n}-\phi_{1}^{*}||l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})|$ (39)
@@
A.3 Almost Sure Convergence of $\phi_{1,n}$
We are now ready to prove Part (a) of Theorem 1 regarding the almost sure convergence of $\phi_{1,n}$. Here and henceforth we will prove more general results by allowing a bias term $\beta_{1,n}=\mathbb{E}\left[\xi_{1,n}\mid\mathcal{G}_{n}\right]$ that may account for errors arising from practical implementations (e.g. the discretization errors; see Remark 2 for details). The following theorem specializes to Part (a) of Theorem 1 when $\beta_{1,n}=\mathbf{0}$.
Theorem 3.
@@ A.3 Almost Sure Convergence of $\phi_{1,n}$
$\mathbb{E}\left[\left|\xi_{1,n}-\beta_{1,n}\right|^{2}\Big{|}\mathcal{G}_{n}\right]\leq$ $cb_{n}(1+|\phi_{1,n}|^{8}+(\log b_{n})^{8})\exp{\{c|\phi_{1,n}|^{6}\}},$ (40)
@@
$(i)$ $\quad 0<a_{n}\leq 1\text{ for all }n,\quad\sum_{n}a_{n}=\infty,\quad\sum_{n}a_{n}|\beta_{1,n}|<\infty;$ (41)
$(ii)$ $\quad c_{1,n}\uparrow\infty,\quad\sum_{n}a_{n}^{2}b_{n}^{q}(\log b_{n})^{q_{2}}e^{cc_{1,n}^{q_{3}}}<\infty$
for any $\quad\text{for any }c>0\text{, }0\leq q\leq 1\text{ , }0\leq q_{2}\leq 8\text{ and }0\leq q_{3}\leq 6;$
@@
Recall that $|\phi_{1,n}|\leq c_{1,n}$ almost surely. By (39), we have
@@
In addition, the assumption (40) yields
@@
$\eta_{1,n}=a_{n}|\beta_{1,n}|+3a_{n}^{2}\bigg{(}$ $c(1+c_{1,n}^{2})e^{cc_{1,n}^{2}}+|\beta_{1,n}|^{2}$ (42)
@@
By assumptions (i)-(ii) of (41), we know $\sum\kappa_{1,n}<\infty$ and $\sum\eta_{1,n}<\infty$. It then follows from (Robbins and Siegmund,, 1971, Theorem 1) that $\left|U_{1,n}\right|^{2}$ converges to a finite limit and $\sum\zeta_{1,n}<\infty$ almost surely.
It remains to show $|U_{1,n}|\to 0$ almost surely. It follows from (35) and (38) that
@@
Now, suppose $|U_{1,n}|\nrightarrow 0$ almost surely. Then there exists a set $Z\in\mathcal{F}$ with $\mathbb{P}(Z)>0$ so that for every $\omega\in Z$, there is $\delta(\omega)>0$ such that $|U_{1,n}|=|\phi_{1,n}-\phi_{1}^{*}|\geq\delta(\omega)>0$ for sufficiently large $n$. Thus,
$\sum\zeta_{1,n}\geq\sum 2\bar{c}a_{n}|\phi_{1,n}-\phi_{1}^{*}|^{2}\geq 2\bar{c}\delta(\omega)^{2}\sum a_{n}=\infty.$
@@
Remark 1.
An instance satisfying the assumptions in (41) is $a_{n}=1\wedge\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}}$, where constants $\alpha>0$, $\beta>0$. $b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}},c_{1,n}=1\vee(\log\log n)^{\frac{1}{6}},$ and $\beta_{1,n}=0$. This is because $\sum\frac{1}{n}=\infty$, and $\sum\frac{(\log n)^{p}(\log\log n)^{q}}{n^{r}}<\infty$, for any $p,q>0$ and $r>1$.
A.4 Mean-Squared Error of $\phi_{1,n}-\phi_{1}^{*}$
In this section, we establish the convergence rate of $\phi_{1,n}$ to $\phi_{1}^{*}$ stipulated in part (b) of Theorem 1.
@@ A.4 Mean-Squared Error of $\phi_{1,n}-\phi_{1}^{*}$
Consider the last inequality in (43) and notice that the left hand side is a decreasing function of $n$ and the right hand side is an increasing function of $n$. So to show that this inequality is true for all $n$, it is sufficient to show that it is true when $n=0$, which is
$(\beta+1)^{\frac{3}{4}}-\beta^{\frac{3}{4}}\leq w\alpha^{\frac{3}{4}}\frac{\beta^{\frac{3}{4}}}{(\beta+1)^{\frac{3}{4}}}.$ (44)
@@
where the last inequality is due to the mean value theorem. Now the desired inequality (44) follows from the fact that $\alpha>\frac{1}{w}$. ∎
The following result covers part (b) of Theorem 1 as a special case.
Theorem 4.
Under the assumption of Theorem 3, if the sequence $\{a_{n}\}$ further satisfies
@@
In particular, if we set the parameters $a_{n},b_{n},c_{1,n},\beta_{1,n}$ as in Remark 1, then
@@
Denote $n_{0}=\inf\{n\geq 0:\phi_{1}^{*}\in K_{1,n+1}\}$ and $U_{1,n}=\phi_{1,n}-\phi_{1}^{*}$. It follows from (35) and (38) that
$\langle U_{1,n},h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})\rangle=-U_{1,n}^{\top}l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})U_{1,n}\leq-\bar{c}|U_{1,n}|^{2}.$ (45)
When $n\geq n_{0}$, this together with a similar argument to the proof of Theorem 3 yields
$\mathbb{E}\left[|U_{1,n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]$ (46)
@@
Moreover, the assumptions in (41) imply that $(1+\frac{1}{3\bar{c}a_{n}})|\beta_{1,n}|^{2}\leq c\frac{b_{n}}{a_{n}}|\beta_{1,n}|^{2}$ for some constant $c>0$. When $n\geq n_{0}$, it follows from (46) that
@@
$c^{\prime}=max\{\frac{\rho_{1}}{a_{0}\hat{\eta}_{1,0}},\frac{\rho_{2}}{a_{1}\hat{\eta}_{1,1}},\cdots,\frac{\rho_{n_{0}+1}}{a_{n_{0}}\hat{\eta}_{1,{n_{0}}}},\frac{3}{\bar{c}}\}+1$. Indeed, it is true when $n\leq n_{0}$. Assume that $\rho_{k+1}\leq c^{\prime}a_{k}\hat{\eta}_{1,k}$ is true for $n_{0}\leq k\leq n-1$. Then (48) yields
@@
In particular, under the setting of Remark 1, we can verify that $(\frac{b_{n}}{a_{n}})|\beta_{1,n}|^{2}$ is a non-decreasing sequence of $n$, and $a_{n}=\Theta(n^{-\frac{3}{4}})$. Then
$\hat{\eta}_{1,n}=$ $cb_{n}(1+c_{1,n}^{8}+(\log b_{n})^{8}+\frac{|\beta_{1,n}|^{2}}{a_{n}})\exp{\{cc_{1,n}^{6}\}}$ (49)
@@
where $c$ and $p$ are positive constants. The proof is now complete.
Appendix B A Proof of Theorem 2
This section is dedicated to proving Theorem 2, which pivots around analyzing the value function $\bar{J}$ in terms of $\phi_{1}$ and $\phi_{2}$.
B.1 Analyzing $\bar{J}(\phi_{1},\phi_{2})$
@@
where $a(\phi_{1})$ is defined by (28) and the functions $f$ and $g$ are defined as follows:
$f(a)=\begin{cases}\frac{x_{0}^{2}(-H-QT)}{2}&\text{if }a=0,\\ \frac{1}{2a}(Q-e^{aT}Q-He^{aT}a)x_{0}^{2}&\text{if }a\neq 0,\end{cases}$ (50)
$g(a)=\begin{cases}\frac{T(-2H-QT)}{4}&\text{if }a=0,\\ \frac{1}{2a^{2}}(QTa+Q+Ha-e^{aT}Q-He^{aT}a)&\text{if }a\neq 0.\end{cases}$ (51)
@@
$\left\{\begin{array}[]{l}v_{t}+\frac{1}{2}\int_{\mathcal{R}^{l}}\sum_{j=1}^{m}(C_{j}x+D_{j}^{\top}u)^{2}\mathcal{N}(u|\phi_{1}x,\phi_{2})\mathrm{d}uv_{xx}\\ \qquad+\left(Ax+B^{\top}\int_{\mathcal{R}^{l}}u\mathcal{N}(u|\phi_{1}x,\phi_{2})\mathrm{d}u\right)v_{x}-\frac{1}{2}Qx^{2}=0,\\ v(T,x)=-\frac{1}{2}Hx^{2}.\end{array}\right.$ (52)
@@
$J(t,x;\phi_{1},\phi_{2})=$ $\frac{1}{2}\left[\frac{Q}{a(\phi_{1})}-e^{a(\phi_{1})(T-t)}(\frac{Q}{a(\phi_{1})}+H)\right]x^{2}-\frac{1}{2}(\sum_{j=1}^{m}D_{j}^{\top}\phi_{2}D_{j})$ (53)
@@
Lemma 5.
Both the functions $f$ and $g$ defined respectively by (50) and (51) are continuously differentiable, monotonically non-increasing, and strictly negative everywhere.
@@
where the inequality is due to the facts that $1+xe^{x}-e^{x}\leq 0$ $\forall x$ and that $Q,H\geq 0$. Moreover, again by L’Hôpital’s rule we obtain

### v4

@@ B.2 Mean Increment
$h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ in the updating rule (4.2). First, note that $\int_{0}^{\cdot}Z_{1,n}^{(2,j)}(t)\mathrm{d}W_{n}^{(j)}(t)$ is a martingale by virtue of Lemma B.1 and (B.6). Taking integration and expectation in (B.5), we get
$\mathbb{E}[Z_{1,n}(s)]$ $=-\int_{0}^{s}k_{1}(t;\boldsymbol{\theta}_{n})(B+(\sum_{j=1}^{m}C_{j}D_{j})+(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\phi_{1,n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t,$ (B.7)
@@
$h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ $=-(B+(\sum_{j=1}^{m}C_{j}D_{j})+(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\phi_{1,n})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t$ (B.8)
@@
$l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})=(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t.$ (B.9)
Next we study $h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$. It follows from (B.1) that $a(\phi_{1})$ is a quadratic function of $\phi_{1}$ and
@@
Hence, by (B.3), we have
$\mathbb{E}[x_{n}(t)^{2}]$ $=(\sum_{j=1}^{m}D_{j}^{\top}\phi_{2,n}D_{j})\int_{0}^{t}e^{a(\phi_{1,n})(t-s)}\mathrm{d}s+x_{0}^{2}e^{a(\phi_{1,n})t}$ (B.10)
@@
$l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})$ $=(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t$ (B.11)
@@
$|h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})|$ $\leq|\phi_{1,n}-\phi_{1}^{*}||l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})|$ (B.12)
@@
B.3 Almost Sure Convergence of $\phi_{1,n}$
We are now ready to prove Part (a) of Theorem 4.1 regarding the almost sure convergence of $\phi_{1,n}$. Here and henceforth we will prove more general results by allowing a bias term $\beta_{1,n}=\mathbb{E}\left[\xi_{1,n}\mid\mathcal{G}_{n}\right]$ that may account for errors arising from practical implementations (e.g. the discretization errors; see the proof of Theorem 4.1 for details). The following theorem specializes to Part (a) of Theorem 4.1 when $\beta_{1,n}=\mathbf{0}$.
Theorem B.3.
@@ B.3 Almost Sure Convergence of $\phi_{1,n}$
$\mathbb{E}\left[\left|\xi_{1,n}-\beta_{1,n}\right|^{2}\Big{|}\mathcal{G}_{n}\right]\leq$ $cb_{n}(1+|\phi_{1,n}|^{8}+(\log b_{n})^{8})\exp{\{c|\phi_{1,n}|^{6}\}},$ (B.13)
@@
$(i)$ $\quad 0<a_{n}\leq 1\text{ for all }n,\quad a_{n}\downarrow 0,\quad\sum_{n}a_{n}=\infty,\quad\sum_{n}a_{n}|\beta_{1,n}|<\infty;$ (B.14)
$(ii)$ $\quad c_{1,n}\uparrow\infty,\quad\sum_{n}a_{n}^{2}b_{n}^{q_{1}}(\log b_{n})^{q_{2}}c_{1,n}^{q_{3}}e^{cc_{1,n}^{q_{4}}}<\infty$
for any $\quad\text{for any }c>0\text{, }0\leq q_{1}\leq 1\text{ , }0\leq q_{2}\leq 8\text{ , }0\leq q_{3}\leq 8\text{ and }0\leq q_{4}\leq 6;$
@@
Recall that $|\phi_{1,n}|\leq c_{1,n}$ almost surely. By (B.12), we have
@@
In addition, the assumption (B.13) yields
@@
$\eta_{1,n}=a_{n}|\beta_{1,n}|+3a_{n}^{2}\bigg{(}$ $c(1+c_{1,n}^{2})e^{cc_{1,n}^{2}}+|\beta_{1,n}|^{2}$ (B.15)
@@
By assumptions (i)-(ii) of (B.14), we know $\sum\kappa_{1,n}<\infty$ and $\sum\eta_{1,n}<\infty$. It then follows from (Robbins and Siegmund, , 1971, Theorem 1) that $\left|U_{1,n}\right|^{2}$ converges to a finite limit and $\sum\zeta_{1,n}<\infty$ almost surely.
It remains to show $|U_{1,n}|\to 0$ almost surely. It follows from (B.8) and (B.11) that
@@
Now, suppose $|U_{1,n}|^{2}\rightarrow c$ almost surely, where $c$ is not almost surely 0. Then, there exists a measurable set $Z\in\mathcal{F}$ with $\mathbb{P}(Z)>0$ such that, for every $\omega\in Z$, there exists a constant $0<\delta(\omega)<c(\omega)$ satisfying
$|U_{1,n}(\omega)|^{2}=|\phi_{1,n}(\omega)-\phi_{1}^{*}|^{2}\geq c(\omega)-\delta(\omega)>0\quad\text{for sufficiently large }n.$
Thus,
$\sum\zeta_{1,n}(\omega)\geq\sum 2\bar{c}a_{n}|\phi_{1,n}(\omega)-\phi_{1}^{*}|^{2}\geq 2\bar{c}(c(\omega)-\delta(\omega))\sum a_{n}=\infty.$
@@
Remark B.4.
An instance satisfying the assumptions in (B.14) is $a_{n}=1\wedge\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}}$, where constants $\alpha>0$, $\beta>0$. $b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}},c_{1,n}=1\vee(\log\log n)^{\frac{1}{6}},$ and $\beta_{1,n}=0$. This is because $\sum\frac{1}{n}=\infty$, and $\sum\frac{(\log n)^{p}(\log\log n)^{q}}{n^{r}}<\infty$, for any $p,q>0$ and $r>1$.
B.4 Mean-Squared Error of $\phi_{1,n}-\phi_{1}^{*}$
In this section, we establish the convergence rate of $\phi_{1,n}$ to $\phi_{1}^{*}$ stipulated in part (b) of Theorem 4.1.
@@ B.4 Mean-Squared Error of $\phi_{1,n}-\phi_{1}^{*}$
Consider the last inequality in (B.16) and notice that the left hand side is a decreasing function of $n$ and the right hand side is an increasing function of $n$. So to show that this inequality is true for all $n$, it is sufficient to show that it is true when $n=0$, which is
$(\beta+1)^{\frac{3}{4}}-\beta^{\frac{3}{4}}\leq w\alpha^{\frac{3}{4}}\frac{\beta^{\frac{3}{4}}}{(\beta+1)^{\frac{3}{4}}}.$ (B.17)
@@
where the last inequality is due to the mean value theorem. Now the desired inequality (B.17) follows from the fact that $\alpha>\frac{1}{w}$. ∎
The following result covers part (b) of Theorem 4.1 as a special case.
Theorem B.6.
Under the assumption of Theorem B.3, if the sequence $\{a_{n}\}$ further satisfies
@@
In particular, if we set the parameters $a_{n},b_{n},c_{1,n},\beta_{1,n}$ as in Remark B.4, then
@@
Denote $n_{0}=\inf\{n\geq 0:\phi_{1}^{*}\in K_{1,n+1}\}$ and $U_{1,n}=\phi_{1,n}-\phi_{1}^{*}$. It follows from (B.8) and (B.11) that
$\langle U_{1,n},h_{1}(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})\rangle=-U_{1,n}^{\top}l(\phi_{1,n},\phi_{2,n};\boldsymbol{\theta}_{n})U_{1,n}\leq-\bar{c}|U_{1,n}|^{2}.$ (B.18)
When $n\geq n_{0}$, this together with a similar argument to the proof of Theorem B.3 yields
$\mathbb{E}\left[|U_{1,n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]$ (B.19)
@@
Moreover, the assumptions in (B.14) imply that $(1+\frac{1}{3\bar{c}a_{n}})|\beta_{1,n}|^{2}\leq c\frac{b_{n}}{a_{n}}|\beta_{1,n}|^{2}$ for some constant $c>0$. When $n\geq n_{0}$, it follows from (B.19) that
@@
$c^{\prime}=max\{\frac{\rho_{1}}{a_{0}\hat{\eta}_{1,0}},\frac{\rho_{2}}{a_{1}\hat{\eta}_{1,1}},\cdots,\frac{\rho_{n_{0}+1}}{a_{n_{0}}\hat{\eta}_{1,{n_{0}}}},\frac{3}{\bar{c}}\}+1$. Indeed, it is true when $n\leq n_{0}$. Assume that $\rho_{k+1}\leq c^{\prime}a_{k}\hat{\eta}_{1,k}$ is true for $n_{0}\leq k\leq n-1$. Then (B.21) yields
@@
In particular, under the setting of Remark B.4, we can verify that $(\frac{b_{n}}{a_{n}})|\beta_{1,n}|^{2}$ is a non-decreasing sequence of $n$, and $a_{n}=\Theta(n^{-\frac{3}{4}})$. Then
$\hat{\eta}_{1,n}=$ $cb_{n}(1+c_{1,n}^{8}+(\log b_{n})^{8}+\frac{|\beta_{1,n}|^{2}}{a_{n}})\exp{\{cc_{1,n}^{6}\}}$ (B.22)
@@
where $c$ and $p$ are positive constants. The proof is now complete. ∎
B.5 Analyzing $\bar{J}(\phi_{1},\phi_{2})$
@@
where $a(\phi_{1})$ is defined by (B.1) and the functions $f$ and $g$ are defined as follows:
$f(a)=\begin{cases}\frac{x_{0}^{2}(-H-QT)}{2}&\text{if }a=0,\\ \frac{1}{2a}(Q-e^{aT}Q-He^{aT}a)x_{0}^{2}&\text{if }a\neq 0,\end{cases}$ (B.23)
$g(a)=\begin{cases}\frac{T(-2H-QT)}{4}&\text{if }a=0,\\ \frac{1}{2a^{2}}(QTa+Q+Ha-e^{aT}Q-He^{aT}a)&\text{if }a\neq 0.\end{cases}$ (B.24)
@@
$\left\{\begin{array}[]{l}v_{t}+\frac{1}{2}\int_{\mathcal{R}^{l}}\sum_{j=1}^{m}(C_{j}x+D_{j}^{\top}u)^{2}\mathcal{N}(u|\phi_{1}x,\phi_{2})\mathrm{d}uv_{xx}\\ \qquad+\left(Ax+B^{\top}\int_{\mathcal{R}^{l}}u\mathcal{N}(u|\phi_{1}x,\phi_{2})\mathrm{d}u\right)v_{x}-\frac{1}{2}Qx^{2}=0,\\ v(T,x)=-\frac{1}{2}Hx^{2}.\end{array}\right.$ (B.25)
@@
$J(t,x;\phi_{1},\phi_{2})=$ $\frac{1}{2}\left[\frac{Q}{a(\phi_{1})}-e^{a(\phi_{1})(T-t)}(\frac{Q}{a(\phi_{1})}+H)\right]x^{2}-\frac{1}{2}(\sum_{j=1}^{m}D_{j}^{\top}\phi_{2}D_{j})$ (B.26)
@@
Lemma B.8.
Both the functions $f$ and $g$ defined respectively by (B.23) and (B.24) are continuously differentiable, monotonically non-increasing, and strictly negative everywhere.
@@
where the inequality is due to the facts that $1+xe^{x}-e^{x}\geq 0$ $\forall x$ and that $Q,H\geq 0$. Moreover, again by L’Hôpital’s rule we obtain